OpenStudy (anonymous):

Ellen wishes to mix candy worth $1.10 per pound with candy worth $2.57 per pound to form 28 pounds of a mixture worth $2.04 per pound. How many pounds of the more expensive candy should she use?

OpenStudy (anonymous):

let x (pounds)=amount of candy worth $1.10 y=amount of candy worth $2.57 then you can build system of equations: x+y=28 1.1*x+2.57*y=2.04*(x+y) please solve it we'll compare answers

OpenStudy (amistre64):

b at $ = t(z-x) ----- ; where t is total pounds; z=final price; x=other price ($-x) and $ = price of amount in question... maybe

OpenStudy (amistre64):

b@2.57 = 28(2.04 - 1.10) -------------- =abt 18lbs ... (2.57 - 1.10)

OpenStudy (amistre64):

17.905 is more accurate

OpenStudy (anonymous):

got the same answer from the system above. i guess, it's the matter of convenience- there are few ways to solve the problem,.... as long as you understand what are you doing... the answer should be the same.

OpenStudy (amistre64):

i never understand what im doing ;)

OpenStudy (amistre64):

a+x = t; a = t-b ax+by=tz (t-b)x +by = tz tx - bx +by = tz b(y-x) = t(z-x) b = t(z-x)/(y-x) ...

OpenStudy (anonymous):

may be sometime... :)) like when you having lunch?

OpenStudy (amistre64):

a+b = t that is lol